# ArguMentor: Augmenting User Experiences with Counter-Perspectives

### v2

@@ Abstract
Opinion pieces (or op-eds) can provide valuable perspectives, but they often represent only one side of a story, which can make readers susceptible to confirmation bias and echo chambers. Exposure to different perspectives can help readers overcome these obstacles and form more robust, nuanced views on important societal issues. We designed ArguMentor, a human-AI collaboration system that highlights claims in opinion pieces, identifies counter-arguments for them using a LLM, and generates a context-based summary of based on current events. It further enhances user understanding through additional features like a Q&A bot (that answers user questions pertaining to the text), DebateMe (an agent that users can argue any side of the piece with) and highlighting (where users can highlight a word or passage to get its definition or context). Our evaluation shows that participants can generate more arguments and counter-arguments and have, on average, have more moderate views after engaging with the system.
@@ Introduction
Opinion pieces form a significant part of our lives. An opinion piece is an article where the writer expresses their opinion on political, social, and societal issues. These often have one characteristic in common: they only represent one side of any story. Coppock, Ekins, and Kirby (2018) find that op-eds are persuasive to both the general public and elites, influencing opinions across society on major issues. Druckman (2005) shows that op-eds can inform major election decisions amongst voters as well. Considering the widespread popularity and significant impact of opinion pieces in our lives, it becomes crucial to address the potential challenges that may emerge while engaging with such content.
@@
Second, online information, such as news articles and posts are often targeted to exploit biases and create divides, driving more clicks and furthering specific agendas (InformationWars). As a result, it becomes nearly impossible to consume unbiased content, as most information is polarized. These factors can lead to the formation of echo chambers within society, driving conversations away from consensus-building discussions.
@@
To combat both idle reading and media bias, researchers have proposed several paper support tools. Idle reading support tools (Fok et al. 2022; August et al. 2023; Kang et al. 2023) support highlighting, critical thinking, question answering, etc for increasing user attention span and understanding of the text. Several papers and websites have tackled media bias by providing multiple perspectives (Park et al. 2009; Perez et al. 2020; Hamborg et al. 2020), guiding users to the same article written from the other side (AllSides Technologies 2012; Ground News Technologies 2024), etc. However, these approaches require the user to take the additional effort to read another article, which users don’t always have the time or motivation for. Moreover, there is a delayed effect, where the user could already be influenced by one article before reading the next one. We argue that users benefit from seeing counter-arguments next to the arguments in real time as they are reading an article.
@@
To address these issues, we present ArguMentor, a human-AI collaboration system that enriches the reading experience of opinion pieces by highlighting claims and generating counter-arguments to help users formulate more robust, nuanced views on important societal issues. ArguMentor provides two types of support: passive and active. The system provides passive support by automatically highlighting main claims in the original text, generating counter-arguments for them, and providing a context based summary for the overall text. ArguMentor provides active support by allowing users to interact with the system by accessing the Q&A bot (that answers user questions pertaining to the text), a DebateMe feature (an agent that users can argue any side of the piece with), and a highlighting trigger window (where users can highlight a word or passage for its definition or context). We derived these features from a preliminary study of news readers and debaters, which illuminated key challenges and potential features and informed our design goals. We evaluated ArguMentor via a within-subjects experiment (N=24). We found that, compared to the baseline, ArguMentor was able to help participants generate more claims, boost their memory for those claims, and even changed participants’ opinions. In summary, the contributions of this paper are as follows:
@@
  • We propose an opinion article reading tool ArguMentor that augments user experience with AI-generated counter-arguments, a debate bot, and a question answering bot.
@@
  • We provide insights and design considerations for future similar tools.
@@
HCI researchers have proposed a variety of software tools and technologies to support reading comprehension and critique of academic papers. To support reading comprehension, Kim et al. (2018) creates an interactive document reader that links the text with its corresponding table cells automatically, which can reduce split attention and facilitate reading. August et al. (2023) created Paper Plain, a system that utilizes NLP techniques to enhance understanding of medical papers. Fok et al. (2022) created a skimming tool that highlights specific parts of the text in different colors to guide the reader’s attention, and enable them to read more efficiently. (Kang et al. 2023) makes the process of finding other papers related to the current paper the user is reading easy by creating a citation graph and threading and summarizing their content using GPT-4. Other recent work (Rachatasumrit et al. 2022; Kang et al. 2022) adds on to this process by automatically highlights important citations, and provides commentary for them based on these papers, and creates a social network for various papers.
To support close reading and critique of written work, Tan et al. (2016) propose a web-based collaborative platform that supports peer interactions and provides feedback for both students and teachers to engage in critical paper reading together. Crebot (Peng et al. 2022) asks users critical questions as they are reading the passage to further their understanding of it. Yuan et al. (2023) uses text summarization techniques and template based questions to help users raise critical thoughts.
@@
Extensive research has been conducted to enhance news consumption through innovative technologies and interfaces.
Laban et al. (2023) design and evaluate news reading interfaces that incorporate discord questions to reveal coverage diversity. Chen et al. (2023) developed Marvista that employs Natural Language Processing (NLP) techniques like abstractive summarization to provide text-specific assistance when users are reading online articles. Their main evaluation study showed that Marvista helps users better comprehend the article. Nguyen et al. (2018) blends information retrieval and human knowledge to create a fact checking portal that aids human fact checkers.
@@
NewsCube (Park et al. 2009) provides readers with multiple perspectives on the same news, by showing them the article written by multiple sides. AllSides Technologies (2012) also does a similar thing by ranking media outlets, and showing a right, left and centrist perspective on the news. Perez et al. (2020) propose a browser extension that presents different perspectives by recommending articles relevant to the current topic. Munson and Resnick (2010) evaluate the extent to which highlighting user agreeable terms within text or showing them first has an impact on their opinion. Hamborg et al. (2020) evaluates word choice bias in media by highlighting trigger words like “terrorists” etc and indicating their positive or negative sentiment.
@@
The use of Large Language Models (LLMs) for tasks related to persuasiveness has been explored in several other contexts. Hyben et al. (2023) tests LLMs and fine-tuned models for claim detection to tackle things like misinformation and spread of bias. Khan et al. (2024) shows that debating with persuasive LLMs leads to truthful answers and shows that debate with an LLM is a good way to resolve conflict in cases where ground truth is unavailable. Breum et al. (2023) shows that chatbots and LLM agents can generate powerful and persuasive arguments, and how they can play an important role in online discussions. Argyle et al. (2023) uses chatbots to show that online political conversations can be improved with an AI assistant’s suggestions. Karinshak et al. (2023) demonstrate that AI can be used to create effective public health messages, and that people are often persuaded by messages created by LLMs.
@@
To brainstorm the potential features a reading experience aiding system could have, we reviewed the aforementioned recent literature in paper reading tools (e.g., (Fok et al. 2022; Kang et al. 2022)), along with asking users for their feedback on a potential list of features. The final list of potential features that is presented to participants of the survey is shown in Table 1.
@@
We recruited 21 participants (primarily young, educated, English-speaking news readers) to complete our survey. The questions focused on their demographics and news reading habits, information about their interaction with opinion pieces, and design ideas for potential technological support. Details of the survey protocol and participant backgrounds can be found in the Supplementary Materials.
@@
  • C1: Difficulty Focusing and Completing the Articles: A majority of the participants indicated that they struggle with long opinion pieces because it is difficult to stay engaged in a long piece. P3 writes, “Opinion pieces are often really long, and I find myself reading a passage, dozing off, then having to re-read it.”
  • C2: Confirmation Bias: Consistent with prior work, a few participants note that they struggle with confirmation bias, where they much prefer reading articles that support their point of view, and avoid reading articles from the other side altogether. P1 notes, “Sometimes its difficult to avoid the tunnel vision that you experience when arguing for or against a topic in a specific way.”
  • C3: Difficulty Imagining Counter-arguments: A few of our participants also noted that they have a hard time thinking about what the other side might say, when they are invested in reading an opinion piece. They indicate that they often have strong opinions about the current piece, one way or another, and can’t easily dismiss those thoughts. When surveyed on whether they go back to researching the other side once they’re done reading, a majority of these participants indicated that they don’t, and often the opinion from the article sticks with them. P1 responded, ”It becomes difficult to think of creative ways to defend/oppose the argument from the templated way in which you are used to doing so.”
@@
Figure 1: Architecture Diagram
Figure 2: ArguMentor System Screenshot
Based on initial feedback and related research, we developed the final architecture for the system, as shown in Figure 1 The system is divided into 2 stages: passive interaction and active interaction. Informed by the preliminary study and our literature review, we created ArguMentor, a human-AI collaboration system that enriches the reading experience of opinion pieces. In this section, we present a user scenario, followed by a description of the back-end architecture.
@@
ArguMentor is deployed on Vercel (Inc. 2015), and Mouseflow (Inc. 2011) is used on the website to track user activity and record sessions.
@@
There are three forms of passive reading support: context of the article, highlighted main claims in the article, and counter arguments for those claims.
@@
    Counter-argument generation: This is a basic instance of in-context few-shot learning (Brown et al. 2020). A few examples are provided in the prompt to show an instance of GPT 3.5 what a potential counter-argument could look like, and it is asked to generate counter-arguments for the claims that are passed to it in a list. This is then displayed on the right hand side of the page. The users can click on ”Expand” to see the whole counter-argument if they are interested in it. Although most times the contexts are efficient and refute the claim, sometimes they can be vague, especially if the claim is about a recent event that the LLM is not trained on.
@@
Active support (interaction) comes primarily in three forms: Q&A, DebateMe, and highlighting.
@@
    DebateMe: This is implemented using LangChain’s conversation chain which enables chains of conversation. Similar to DebateDevil, the user can enter their argument, and the bot will debate the other side. This is implemented using prompting — the LLM is asked to debate the other side to the user’s argument, and be as brief and persuasive as possible. The user is then given an option to click on a ”thumbs up” or ”thumbs down” button. If the user clicks on ”thumbs down”, the LLM will find another way to persuade the user. This is done to ensure that the user is satisfied with the response, and can report any problematic content to the developers. This is also done using in-context prompting.
@@
To investigate the benefits of ArguMentor for users reading opinion pieces, we conducted a within-subjects experiment with 24 participants. In this experimental design, each participant read some opinion articles using ArguMentor (experimental condition) and others without the system (control condition), i.e., typical reading without any intervention.
Task and Procedure:
@@
Figure 3: Experiment Design and Procedure
First, participants give their stance on the main topic of six articles on a scale of 1–5 (1=strongly disagree, 5=strongly agree). Then, participants are given a polarizing article from the right in the baseline (no ArguMentor) condition. A polarizing article is defined as one that is based on one of the major talking points of the political right or left (e.g., abortion, immigration), where most people would have a strong reaction to the article one way or another. Although the terms ”left”, ”right” and ”neutral” as used in this paper refer to American politics, the articles don’t require the users to know about American politics to be able to participate. They are then asked to fill a survey which asks for how many claims they can spot, and counter-arguments they can think of. They are also asked to answer some basic attention checking questions. This is repeated for left-leaning articles, and finally a neutral article. A break is given after this procedure.
Then, participants repeat this process for three additional articles (representing left, right, and neutral perspectives) in the experimental (ArguMentor) condition, including the post-surveys. TA final, exit survey consisting of subjective experience and system feedback questions is given in the end. Participants can have as much time with each article and the system as they wish. They are instructed to not return to the article or the system while filling the survey, and answer subjective questions as honestly as possible.
The document selection is counterbalanced to reduce order effects; i.e., Participant 1 gets Documents A–C in the control condition and D–F in the experimental condition, Participant 2 gets Documents D–F in the control condition and A–C in the experimental condition, and so on. However, all participants experienced the experimental condition after the control condition.
The articles are chosen from AllSides Technologies (2012), which collates daily top news stories from the political left, right, and center. To create a spectrum of reading comprehension issues, the article content is tweaked using ChatGPT to make one article more challenging to read (adding difficult language, jargon, etc), one medium difficulty, and one easy to read. The six articles are included in the Supplementary Materials.
Evaluation Metrics:
@@
We measured participants’ outcomes in two areas: finding claims and thinking of counter-arguments for those claims. As our preliminary study indicated, users have a hard time doing both of these tasks, so we want to see if the system improves these skills. We count the number of claims/counterarguments participants submit in each post survey. Our hypothesis is that use of ArguMentor will increase the number of both.
@@
If the participants can come up with an equal number of arguments for both, but the system takes them twice as long to use, it might suggest that the system is burdensome to use; whereas the benefits are clearer if the participant can come up with more counter arguments after using the system for more or less an equal time than just reading the article. However, it is important to consider that in contexts where efficiency is not a priority, such as leisure reading, taking more time might be acceptable if the system offers a better subjective experience. We also want to measure how the user spends time on the system– which features they use, for how long, etc.
@@
We measure the participant’s subjective experience of the system on two metrics: how they feel about the system, and how they feel about the article after using the system. Specifically, the survey captures how strongly the participant felt about a topic without the system (on a scale of 1–5), and then how they feel about it after interacting with the system. This way we can test if the counter-arguments and the system has an impact on participants’ perception of the topic. After that, we can adopt the technology acceptance model as well as a survey to take feedback on the system.
Participants:
We recruited 24 participants from our university and an online quizzing group. A majority speak English as their first language, read the news regularly, and come from a STEM background. Half identified as men; the others as women. All have at least a master’s degree in formal education.

### v3

@@ Abstract
We encounter arguments everyday in the form of social media posts, presidential debates, news articles, and even advertisements. A ubiquitous, influential example is the opinion piece (op-ed). Opinion pieces can provide valuable perspectives, but they often represent only one side of a story, which can make readers susceptible to confirmation bias and echo chambers. Exposure to different perspectives can help readers overcome these obstacles and form more robust, nuanced views on important societal issues. We designed ArguMentor, a human-AI collaboration system that highlights claims in opinion pieces, identifies counter-arguments for them using a LLM, and generates a context-based summary of based on current events. It further enhances user understanding through additional features like a Q&A bot (that answers user questions pertaining to the text), DebateMe (an agent that users can argue any side of the piece with) and highlighting (where users can highlight a word or passage to get its definition or context). Our evaluation on news op-eds shows that participants can generate more arguments and counter-arguments and display higher critical thinking skills after engaging with the system. Further discussion highlights a more general need for this kind of a system.
@@ Introduction
People encounter and process arguments in a variety of forms every day. Argumentation is a core part of daily decision-making, as we constantly weigh different perspectives and challenge assumptions to reach well-informed conclusions. Engaging in debate and resolving it allows us to refine our thoughts, test the strength of our reasoning, and understand diverse viewpoints. This process is essential for making better decisions, as it encourages critical thinking and helps us navigate complex issues.
Opinion pieces in the news are one form of this type of debate. An opinion piece (or op-ed) is an article where the writer expresses their opinion on political, social, and societal issues. These op-eds often have one characteristic in common: they only represent one side of any story. Coppock et al. (Coppock, Ekins, and Kirby 2018) find that op-eds are persuasive to both the general public and elites, influencing opinions across society on major issues. Druckman (Druckman 2005) shows that op-eds can inform major election decisions amongst voters as well. Considering the widespread popularity and significant impact of opinion pieces in our lives, it becomes crucial to address the potential challenges that may emerge while engaging with such content.
@@
Second, online information, such as news articles and social media posts, is often designed to exploit biases and create divides, driving more clicks and furthering specific agendas (Stengel 2019). As a result, it becomes nearly impossible to consume unbiased content, as most information is polarized. These factors can lead to the formation of echo chambers within society, driving conversations away from consensus-building discussions. To combat both idle reading and media bias, researchers have proposed several reading support tools. Idle reading support tools (Fok et al. 2022; August et al. 2023; Kang et al. 2023) support highlighting, critical thinking, question answering, etc for increasing user attention span and understanding of the text. Several research projects and commercial websites have tackled media bias by providing multiple perspectives (Park et al. 2009; Perez et al. 2020; Hamborg et al. 2020) or guiding users to the same article written from the other side (AllSides Technologies 2012; Ground News Technologies 2024). However, these approaches require the user to take the additional effort to read another article, which users don’t always have the time or motivation for. Moreover, there is a delayed effect, where the user could already be influenced by one article before reading the next one. (Brashier et al. 2021) shows that timing is often the most critical part of fact checking. Hence, we argue that users benefit from seeing counter-arguments next to the arguments in real time as they are reading an article.
@@
To address these issues, we present ArguMentor, [^1] We will release the ArguMentor system as open source software and the articles used if this paper is accepted. a human-AI collaboration system that enriches the reading experience of opinion pieces by highlighting claims and generating counter-arguments to help users formulate more robust, nuanced views on important societal issues. ArguMentor provides two types of support: passive and active. The system provides passive support by automatically highlighting main claims in the original text, generating counter-arguments for them, and providing a context based summary for the overall text. ArguMentor provides active support by allowing users to interact with the system by accessing the Q&A bot (that answers user questions pertaining to the text), a DebateMe feature (an agent that users can argue any side of the piece with), and a highlighting trigger window (where users can highlight a word or passage for its definition or context). We derived these features from a preliminary study of news readers, which illuminated key challenges and potential features and informed our design goals.
We evaluated ArguMentor via a within-subjects experiment (N=24). We found that, compared to the baseline, ArguMentor was able to help participants generate more claims and counter-arguments, boost their memory for those claims, and even gain better critical thinking skills. Follow-up interviews showed how this system aids those with lesser vocabulary or reading skills, where reading an article is more challenging. In summary, the contributions of this paper are as follows:
@@
  • We propose an opinion article reading tool, ArguMentor, that augments the user experience with tailored AI-generated counter-arguments, a debate bot, and a question answering bot.
@@
  • We provide insights and design considerations for future similar tools to enhance critical thinking.
@@
HCI researchers have proposed a variety of software tools and technologies to support reading comprehension and critique of academic papers. To support reading comprehension, Kim et al. (Kim et al. 2018) created an interactive document reader that links the text with its corresponding table cells automatically, which can reduce split attention and facilitate reading. August et al. (August et al. 2023) created Paper Plain, a system that utilizes NLP techniques to enhance understanding of medical papers. Fok et al. (Fok et al. 2022) created a skimming tool that highlights specific parts of the text in different colors to guide the reader’s attention, and enable them to read more efficiently. Kang et al. (Kang et al. 2023) makes the process of finding other papers related to the current paper the user is reading easy by creating a citation graph and threading and summarizing their content using GPT-4. Other recent work (Rachatasumrit et al. 2022; Kang et al. 2022) adds on to this process by automatically highlights important citations, and provides commentary for them based on these papers, and creates a social network for various papers.
To support close reading and critique of written work, (Tan et al. 2016) propose a web-based collaborative platform that supports peer interactions and provides feedback for both students and teachers to engage in critical paper reading together. Crebot (Peng et al. 2022) asks users critical questions as they are reading the passage to further their understanding of it. Critrainer (Yuan et al. 2023) uses text summarization techniques and template based questions to help users raise critical thoughts.
@@
Extensive research has been conducted to enhance news consumption through novel technologies and interfaces. Laban et al. (Laban et al. 2023) design and evaluate news reading interfaces that incorporate discord questions to reveal coverage diversity. Chen et al. (Chen et al. 2023) developed Marvista that employs Natural Language Processing (NLP) techniques like abstractive summarization to provide text-specific assistance when users are reading online articles. Their main evaluation study showed that Marvista helps users better comprehend the article. Nguyen et al. (Nguyen et al. 2018) blends information retrieval and human knowledge to create a fact checking portal that aids human fact checkers.
@@
NewsCube (Park et al. 2009) provides readers with multiple perspectives on the same news, by showing them the article written by multiple sides. AllSides (AllSides Technologies 2012) also does a similar thing by ranking media outlets, and showing a right, left and centrist perspective on the news. Perez et al. (Perez et al. 2020) propose a browser extension that presents different perspectives by recommending articles relevant to the current topic. Munson and Resnick (Munson and Resnick 2010) evaluate the extent to which highlighting user agreeable terms within text or showing them first has an impact on their opinion. Hamborg et al. (Hamborg et al. 2020) evaluate word choice bias in media by highlighting trigger words like “terrorists” etc and indicating their positive or negative sentiment.
@@
The use of Large Language Models (LLMs) for tasks related to persuasiveness has been explored in several other contexts. Hyben et al. (Hyben et al. 2023) tested LLMs and fine-tuned models for claim detection to tackle things like misinformation and spread of bias. Khan et al. (Khan et al. 2024) show that debating with persuasive LLMs leads to truthful answers and show that debate with an LLM is a good way to resolve conflict in cases where ground truth is unavailable. Breum et al. (Breum et al. 2023) show that chatbots and LLM agents can generate powerful and persuasive arguments, and how they can play an important role in online discussions. Argyle et al. (Argyle et al. 2023) uses chatbots to show that online political conversations can be improved with an AI assistant’s suggestions. Karinshak et al. (Karinshak et al. 2023) demonstrate that AI can be used to create effective public health messages, and that people are often persuaded by messages created by LLMs.
@@
To brainstorm the potential features a reading experience aiding system could have, we reviewed the aforementioned recent literature in paper reading tools (e.g., (Fok et al. 2022; Kang et al. 2022)), along with creating a survey asking users for their feedback on a potential list of features. The final list of potential features that is presented to participants of the survey is shown in Table 1.
@@
We recruited 40 participants to complete our survey. The age range was 20–45. Fifty-five percent were men and 45% were women. For 30%, English was their second or higher language. There was a diverse mix of educational backgrounds, from those who did not complete high school, to college graduates, to those with advanced degrees. Around 70% of them were frequent news readers. These participants were selected through snowball sampling participants at our university and an online gaming group (these groups tend to attract a diverse range of individuals from various backgrounds). The questions focused on their demographics and news reading habits, information about their interaction with opinion pieces, and design ideas for potential technological support. Details of the survey protocol and participant backgrounds can be found in the Supplementary Materials.
@@
  • C1: Difficulty Focusing and Completing the Articles: A majority of the participants indicated that they struggle with long opinion pieces because it is difficult to stay engaged in a long piece. For example, P3 wrote, “Opinion pieces are often really long, and I find myself reading a passage, ‘dozing off, then having to re-read it.”
  • C2: Confirmation Bias: Consistent with prior work, a few participants note that they struggle with confirmation bias, where they much prefer reading articles that support their point of view, and avoid reading articles from the other side altogether. For example, P1 noted, “Sometimes its difficult to avoid the tunnel vision that you experience when arguing for or against a topic in a specific way.”
  • C3: Difficulty Imagining Counter-arguments: A few of our participants also noted that they have a hard time thinking about what the other side might say, when they are invested in reading an opinion piece. They indicate that they often have strong opinions about the current piece, one way or another, and can’t easily dismiss those thoughts. When surveyed on whether they go back to researching the other side once they’re done reading, a majority of these participants indicated that they don’t, and often the opinion from the article sticks with them. For example, P1 responded, ”It becomes difficult to think of creative ways to defend/oppose the argument from the templated way in which you are used to doing so.”
@@
Informed by the preliminary study and our literature review, we created ArguMentor, a human-AI collaboration system that enriches the reading experience of opinion pieces. In this section, we present a user scenario, followed by a description of the back-end architecture.
@@
ArguMentor is deployed on Vercel (Inc. 2015), and Mouseflow (Inc. 2011) is used on the website to track user activity and record sessions. The system is divided into 2 stages: passive interaction and active interaction (see Figure 1).
Figure 1: Architecture Diagram
@@
There are three forms of passive reading support: context of the article, highlighted main claims in the article, and counter arguments for those claims (see Figure 3).
@@
    Counter-argument generation: To generate effective counter-arguments, we fine-tune a GPT 3.5 based model using a rebuttal dataset (Orbach et al. 2019). The model was trained on 32,176 tokens across 8 epochs with a batch size of 1 and a learning rate multiplier of 2. The final training loss is shown in Figure 2. LLMs by themselves are shown to generate effective rebuttals, but fine-tuning a model on a specific dataset ensures that the rebuttals are even more effective. (Carrasco-Farre 2024) show that these rebuttals, generated by LLMs are targeted, specific and persuasive, which would achieve a better result than just in-context learning. These counter-arguments are then displayed on the right hand side of the page. The users can click on ”Expand” to see the whole counter-argument if they are interested in it. Although most times the contexts are efficient and refute the claim, sometimes they can be vague, especially if the claim is about a recent event that the LLM is not trained on.
    Figure 2: Training Loss: OpenAI GPT 3.5 training loss by training over the IBM Rebuttals Dataset
@@
Active support (interaction) comes primarily in three forms: Q&A, DebateMe, and highlighting (see Figure 3).
@@
    DebateMe: This is implemented using LangChain’s conversation chain which enables chains of conversation. Similar to (DebateDevil 2022), the user can enter their argument, and the bot will debate the other side. This is implemented using prompting — the LLM is asked to debate the other side to the user’s argument, and be as brief and persuasive as possible. The user is then given an option to click on a ”thumbs up” or ”thumbs down” button. If the user clicks on ”thumbs down”, the LLM will find another way to persuade the user. This is done to ensure that the user is satisfied with the response, and can report any problematic content to the developers. This is also done using in-context prompting.
@@
To investigate the benefits of ArguMentor for users reading opinion pieces, we conducted a within-subjects experiment with 24 participants. In this experimental design, each participant read one op-ed using ArguMentor (experimental condition) and one without the system (control condition), i.e., typical reading without any intervention.
Baseline
We select reading the article without any intervention (no system, or any other support tool) as our baseline. We justify the choice of our baseline here:
  • First, most readers in our preliminary survey indicated that this is how they consumed news on a regular basis. We know this to be representative of the general population as well. In that case, it makes the most sense to select this as the baseline.
  • Secondly, even though websites like GroundNews and AllSides exist, they work very differently to our system. For example, ground news categorizes a news piece as left-, right-, or center-leaning. It then identifies how that issue has been covered by all three sides and various outlets. There are no counter-arguments, other than if the right covered a topic completely different from the left. On the other hand, our system provides direct rebuttals to a given claim. It does not make the users go through another article written from a different viewpoint.
Articles
To ensure a diverse selection of op-ed articles, we chose six articles from the news spanning three subfields: politics, social issues, and economics. For politics, we focused on the U.S. election as a key topic; for economics, we selected stock buybacks; and for social issues, we focused on feminism. This range represents a broad spectrum of socio-economic topics, allowing us to evaluate the effectiveness of our system across different subject areas. Two articles are chosen per key topic, and each article can be read with our system or with the baseline (without the system), creating 12 unique conditions. Participants weren’t required to know about the topics in order to participate in the study.
Task and Procedure
@@
Figure 4: Experiment Design and Procedure: The process always starts with the baseline condition (reading the article by itself), and ends with the experimental condition (reading the article with the system). Each person only reads two articles and submits questionnaires right after reading.
First, demographic information is collected for all participants. Then, participants are given the baseline condition. The baseline condition is always presented first because it avoids participants finding out what the ArguMentor system (experimental condition) looks like and attempting to replicate its functionality later for the baseline.
Next, participants are presented with a survey. The survey asks several questions. First, it asks attention checking questions to see if the participants understood the basic facts in the article. These are multiple choice questions such as what was the topic of article, who did the article support, what is the main point of the article, etc.
The survey then asks participants to write any claims, counter-arguments, questions, biases, and solution to problems they can think of, based on the article. They are then asked to answer ACT-style questions related to deduction, assumptions the author makes, whether the author should remove or keep certain sentences to prove their point, etc. These questions are generated by prompting a Llama-3 model to generate ACT style questions based on a paragraph. The authors have then gone through each question in detail to ensure they are answerable and valid. The claims and counter-arguments are timed, whereas the critical thinking portion is left untimed. This is because critical thinking can take some time, and we give participants the flexibility to take breaks, etc to think of the given questions/prompts. We time claim and counter-argument generation to see the immediate effectiveness of our system. The choice for selecting these metrics is explained in the next section.
This procedure is then repeated with the experimental condition, with participants getting access to ArguMentor. A final, exit survey consisting of subjective experience and system feedback questions is given in the end. Participants can have as much time with each article and the system as they wish. They are instructed to not return to the article or the system while filling the survey, and answer subjective questions as honestly as possible.
The document selection is counterbalanced to reduce order effects; i.e., Participant 1 gets Document A in the control condition and B in the experimental condition, Participant 2 gets Documents B in the control condition and A in the experimental condition, and so on. However, all participants experienced the experimental condition after the control condition. 24 people read 2 articles each, resulting in 48 data points for our 12 conditions (6 articles, each with and without the system).
Evaluation Metrics
@@
We measured participants’ outcomes in two areas: (1) finding claims and (2) thinking of counter-arguments for those claims. As our preliminary study indicated, users have a hard time doing both of these tasks, so we want to see if the system improves these skills. These two skills are also fundamental for everyday tasks; i.e., being able to spot the main claim in a news article, social media post, or a live debate and then thinking of a potential counter to it to see what the other side would say. We count the number of claims / counterarguments participants submit in each post survey. Our hypothesis is that use of ArguMentor will increase the number of both.
We also measure how many biases present in the article users can write about, how many questions they pose, how many solutions they pose to the problem in the article, and how many ACT style questions they answer correctly (out of 10). This is the critical thinking portion of the experiment. These are the facets of critical thinking tests such as (Technologies 2024) as described in (Wyn Davies ). The Watson Glaser Critical Thinking Test measures ”RED” – Recognize Assumptions, Evaluate Arguments, Draw Conclusions through various steps. We want to see if our system erodes or promotes critical thinking. The system only gives them counter-arguments, and the ability to ask questions related to the article, how creatively they use it for these tasks, and whether the ultimate outcome is better are metrics to be evaluated.
@@
If the participants can come up with an equal number of arguments for both, but the system takes them twice as long to use, it might suggest that the system is burdensome to use. The benefits are clearer if the participant can come up with more counter arguments after using the system for more or less an equal time than just reading the article. However, it is important to consider that in contexts where efficiency is not a priority, such as leisure reading, taking more time might be acceptable if the system offers a better subjective experience. We also want to measure how the user spends time on the system: which features they use, for how long, etc.
@@
We measure the participant’s subjective experience of the system by asking them how they feel about the system. We adopt the technology acceptance model as well as a survey to collect feedback on the system. A short interview is also conducted to allow them to express their genuine thoughts.
Quality Analysis
Participants might just generate more arguments with the system, but that doesn’t convey how strong these arguments are. To gauge this, we compare the quality of counter-arguments the participants generate with the baseline. This is done in two ways: computational and human. Increasingly, research has shown the prominence of GPT-as-judge (Zheng et al. 2023). We deploy this method to ask a model the binary question, ”which [solution/argument] is stronger: x (baseline) or y (system)?”. This is compared across topics and participants. To avoid any potential bias such as a model preferring an argument that sounds like its own, we use a Llama-3 7b model for this task.
Secondly, we also recruited three expert debaters, who have judged and ”broken in” (a debate term for qualifying for out rounds such as semi-finals or finals) over 3 international parliamentary debating competitions such as UADC, Australs, etc. We ask them the same binary question about a randomly selected counter-argument from each survey we collect (each expert answered 48 questions). These counter-arguments are presented in random order (from the baseline or the experimental condition). We check their agreement on this task, and only consider the argument as superior if all 3 experts voted for it. In case of a non-unanimous answer, we ignore that data point.
Participants
We initially recruited 5 participants from a gaming convention group. Each was then instructed to recommend another person for the study (snowball sampling) until a total of 24 participants were recruited. Eleven participants were men, 11 were women, and 2 identified as other genders. 33% spoke English as their first language, while a majority spoke at least one other language. Most of them reported reading the news at least semi-regularly (twice a week). The participants ranged from 19–43 years old, representing a vast range of individuals. They also represented a variety of income brackets, with over half being middle or lower-middle class. Ten had only finished high school or had not finished high school, 8 had a bachelor’s degree, and 6 held advanced degrees.